OpenStudy (anonymous):

Simplify (4.5)(−9)(−1.8). −729 −72.9 72.9 729

OpenStudy (anonymous):

Lol i was just about to ask you for help

OpenStudy (michele_laino):

at first step, we can write this: \((4.5 1 \cdot (-9)) \cdot (-1.8)\)

OpenStudy (anonymous):

i know its going to be positive answer

OpenStudy (michele_laino):

yes! and you have to apply the associative property of multiplication

OpenStudy (anonymous):

729

OpenStudy (michele_laino):

so, what is \(4.5 \cdot (-9)=...?\)

OpenStudy (michele_laino):

I think that 729 is wrong

OpenStudy (angelmarie328):

72.9 is the answer.